Parallel connection solve for the total capacitance, total charge and individual charge.

C1 = 6.31 x 10^-12

C2 = 3.75 x 10^-12

C3 = 7.23 x 10^-12

V = 1400 V


1
Expert's answer
2021-10-31T18:13:17-0400

"C=C_1+C_2+C_3=17.29\cdot 10^{-12}~F,"

"q=CU=24.2\cdot 10^{-9}~C,"

"q_1=C_1U=8.8\cdot 10^{-9}~C,"

"q_2=C_2U=5.3\cdot 10^{-9}~C,"

"q_3=C_3U=10.1\cdot 10^{-9}~C."